# Looking at Model Debiasing through the Lens of Anomaly Detection

### v2

@@ 4.3.2 Realistic Datasets
Table 3 reports the performance of MoDAD and a comparison with relevant existing works on the three real-world datasets. First, we evaluate the accuracy of our bias-identification approaches on Waterbirds and BFFHQ, obtaining an average f1-score of $0.70$ and $0.75$, respectively (see Table 2). Regarding the debiasing performance, MoDAD surpasses other unsupervised debiasing works, apart from the Average Accuracy on Waterbirds, where LfF [12] reports the highest among all the works in this category. At the same time, our method reaches higher Conflicting Accuracy, almost matching the supervised state-of-the-art on this dataset (GroupDRO [1]). Regarding BAR, we match the performance of DebiAN [31] ($-0.05\%$) on average, but with a significantly lower standard deviation ($2.92$ for DebiAN, $0.72$ for MoDAD). BFFHQ has a bias-correlation $\rho=0.995$, thus capturing the contribution of the few bias-conflicting training is much harder. However, in BFFHQ we report the most significant improvement over existing works, with a Conflicting Accuracy of $68.33\%$, corresponding to an improvement of $4.46\%$ over the second-best approach (DFA [13]).
@@ 5 Conclusions
In our bias identification step, we customize the OCSVM algorithm, manipulating its decision function to accommodate a custom per-class decision threshold $\tau_{y}$. These thresholds (different for each class) push the anomaly detectors to reject samples as out-of-class (i.e.., bias-conflicting) only when they are highly confident, and thus better discriminating between bias-aligned and bias-conflicting samples. In Table S.1, we report an ablation study on Corrupted CIFAR-10 (for three considered bias correlation values $\rho$), evaluating the impact of using our custom thresholds or exploiting the original formulation with the default threshold (which is 00).
@@
Our method (MoDAD) is designed to be applied to a biased model, i.e.., on a model that fails to generalize even when properly trained and tuned, suggesting the presence of bias in data or dramatic distribution shifts. In such a case, MoDAD has a significant impact, succeeding in reducing the bias effect and debiasing the model. However, in general, we do not know whether a dataset is biased or not. Hence, we are interested in assessing the effect of MoDAD when applied to an unbiased model, i.e.., a model trained on an unbiased dataset. Thus, we run the two steps of MoDAD on the original non-biased version of CIFAR-10 and estimate average accuracy on the test set. As we can see in Table S.2, our debiasing method affects only slightly the performance on the test set with respect to the baseline ERM+CE trained model: we experience just a drop of $\sim 2\%$, which is significantly lower than the average benefit that we can get when using MoDAD on actually biased datasets.

### v3

@@ 4.3.2 Realistic Datasets
Table 3 reports the performance of MoDAD and a comparison with relevant existing works on the three real-world datasets. First, we evaluate the accuracy of our bias-identification approaches on Waterbirds and BFFHQ, obtaining an average f1-score of $0.70$ and $0.75$, respectively (see Table S.2). Regarding the debiasing performance, MoDAD surpasses other unsupervised debiasing works, apart from the Average Accuracy on Waterbirds, where LfF [25] reports the highest among all the works in this category. At the same time, our method reaches higher Conflicting Accuracy, almost matching the supervised state-of-the-art on this dataset (GroupDRO [30]). Regarding BAR, we match the performance of DebiAN [18] ($-0.05\%$) on average, but with a significantly lower standard deviation ($2.92$ for DebiAN, $0.72$ for MoDAD). BFFHQ has a bias-correlation $\rho=0.995$, thus capturing the contribution of the few bias-conflicting training is much harder. However, in BFFHQ we report the most significant improvement over existing works, with a Conflicting Accuracy of $68.33\%$, corresponding to an improvement of $4.46\%$ over the second-best approach (DFA [17]).
@@ 5 Conclusions
Acknowledgments
We acknowledge the financial support from PNRR MUR Project PE0000013 "FAIR", funded by the European Union - NextGeneration EU, CUP J33C24000430007
@@ Acknowledgments
In our bias identification step, we customize the OCSVM algorithm, manipulating its decision function to accommodate a custom per-class decision threshold $\tau_{y}$. These thresholds (different for each class) push the anomaly detectors to reject samples as out-of-class (i.e., bias-conflicting) only when they are highly confident, and thus better discriminating between bias-aligned and bias-conflicting samples. In Table S.1, we report an ablation study on Corrupted CIFAR-10 (for three considered bias correlation values $\rho$), evaluating the impact of using our custom thresholds or exploiting the original formulation with the default threshold (which is 00).
@@
Our method (MoDAD) is designed to be applied to a biased model, i.e., on a model that fails to generalize even when properly trained and tuned, suggesting the presence of bias in data or dramatic distribution shifts. In such a case, MoDAD has a significant impact, succeeding in reducing the bias effect and debiasing the model. However, in general, we do not know whether a dataset is biased or not. Hence, we are interested in assessing the effect of MoDAD when applied to an unbiased model, i.e., a model trained on an unbiased dataset. Thus, we run the two steps of MoDAD on the original non-biased version of CIFAR-10 and estimate average accuracy on the test set. As we can see in Table S.2, our debiasing method affects only slightly the performance on the test set with respect to the baseline ERM+CE trained model: we experience just a drop of $\sim 2\%$, which is significantly lower than the average benefit that we can get when using MoDAD on actually biased datasets.